OpenStudy (anonymous):

Multiply and simplify: x+3/ x divided by x+4/ x^2+7x+12?

OpenStudy (anonymous):

\[\frac{x + 3}{x} \div \frac{x + 4}{x^{2} + 7x + 12}\] \[\frac{x + 3}{x} \times \frac{(x + 3)(x + 4)}{x + 4}\] \[\frac{x + 3}{x} \times \frac{(x + 3)}{1}\] \[\frac{(x + 3)^{2}}{x}\] \[\frac{x^{2} + 6x + 9}{x}\]

OpenStudy (anonymous):

thx so much! i was really struggling with this :)

OpenStudy (anonymous):

Anytime :)